A sample of argon gas that originally occupied 14.6L at 25.0oC was heated to 50.0oC at constant pressure. What is its new volume?
1
Expert's answer
2020-09-15T13:52:32-0400

Charles law formula is


"\dfrac{V1}{T1}=\dfrac{V2}{T2}" Where,

V1=14.6L

V2=??

Converting temperature into kelvin

T1=25+273=298K

T2=50+273=323k

Substituting values into the formula

"\dfrac{14.6}{298}=\dfrac{v2}{323}"

Making V2 the subject

"V2=\dfrac{14.6*323}{298}"


"V2=15.8L"

The new volume is "15.8L"